# Davi Lazzarotto, Michela Testolina, Touradj Ebrahimi Multimedia Signal Processing Group (MMSPG) École Polytechnique Fédérale de Lausanne (EPFL) Lausanne, Switzerland

### v1

@@ III-B Objective quality metrics
III-C Benchmarking settings
Prior to the assessment of the quality predictors, the objective scores were fitted to the MOS following the procedure reported in [44], adopting the logistic function with offset depicted in Equation 1.
$S_{p}=\beta_{1}+\frac{\beta_{2}}{1+\exp(-\beta_{3}*(S-\beta_{4}))}$ (1)
The performance of the objective quality metrics was then evaluated through multiple performance indexes, including the correlation between subjective and fitted objective scores using both the Pearson Linear Correlation Coefficient (PLCC) and the Spearman’s Rank Correlation Coefficient (SROCC). In addition, the Root Mean Square Error (RMSE) and the Outlier Ratio (OR) were also computed following the guidelines in ITU-T Recommendation P.1401 [45].
IV Results and discussion
The obtained performance indexes for the entire dataset are displayed in the first four columns of Table I. None of the analyzed metrics is able to achieve PLCC higher than 0.9 or SROCC higher than 0.8, suggesting that current metrics are still not able to fully predict the human subjects quality assessment in the 3D domain. The image-based FSIM metric achieves the best performance index values, followed by VMAF, PCQM, and MS-PointSSIM. The advantage of using image quality metrics in this case may be because they consider the information on how the point clouds were rendered. In fact, changing the point size on the subjective assessment platform could have an impact on the subjective scores, which would be reflected in image metrics while point cloud metrics would remain unchanged. Apart from FSIM and VMAF, MS-SSIM and IW-SSIM are also able to achieve good performance. Regarding point cloud metrics, PCQM and MS-PointSSIM achieve the highest performances, the first with an advantage on SROCC and the latter with lower OR. GraphSIM and MS-GraphSIM come next, displaying better performance than the two best geometry-based metrics, i.e. MPED and D2 PSNR. However, while the multi-scale paradigm allows MS-PointSSIM to outperform PointSSIM by a large margin, it does not have the same effect on GraphSIM, which achieves a slightly higher performance than MS-GraphSIM. As expected, RR and NR quality metrics display lower performance than the evaluated FR metrics. Within those categories, the use of multi-modal learning allowed MM-PCQA to outperform VQA-PC, which relies only on projections onto a moving camera. The use of 2D saliency maps on RR-CAP bring superiority when compared to the 3D-based features used by PCM_RR. Figure 2 displays the scatter plots of the objective scores against the MOS, further illustrating the predictive ability of the best-performing metrics as well as other cases of interest.
@@ IV Results and discussion
Although the performance indexes computed across the entire dataset provide a good indication of both cross-content and cross-codec generalization, there are scenarios where requirements may be lower. For instance, when a metric is used to estimate the quality of point clouds compressed with a known codec, only cross-content generalization could be desired, which can be better evaluated by the six right-most columns of Table I. It is observed that correlation values with G-PCC are usually higher than for the other codecs, with many metrics achieving PLCC higher than 0.9 and significantly better performance than for the entire dataset. A possible explanation is the simplicity of octree-based geometric artifacts, as stronger compression always leads to fewer points in the decoded point cloud and sparser distribution. GraphSIM achieves the best performance in this case, followed by MS-GraphSIM and image-based VMAF. However, these metrics were not able to maintain high performance for V-PCC, where sharp drops in performance are observed for both point-to-distribution variants, Histogram Y and YUV, GraphSIM, MS-GraphSIM, and PointSSIM. For the last three metrics, the reason can be observed in Figure 2, where stimuli generated from Boxer and House_without_roof with low objective scores received a high MOS. These two models have different behavior likely due to their lower point density. Image-based FSIM and PCQM achieve the highest performance, followed by MS-PointSSIM in terms of PLCC and MM-PCQA according to the SROCC. For JPEG Pleno, the correlation coefficients are in between G-PCC and V-PCC, with GraphSIM and LogP2D-JGY achieving the highest correlation values.
@@
Some applications prioritize cross-codec generalization, especially when comparing the performance between compression standards. Since rate-distortion curves are usually compared separately for each point cloud, cross-content generalization is less important. Table II illustrates the PLCC and SROCC computed separately for each point cloud, where most objective metrics achieve higher correlation when compared to previous evaluations. The highest coefficients are achieved for Bouquet, Soldier, and Thaidancer, which are solid point clouds, i.e. with very high point density. The higher voxelization precision of the latter does not seem to negatively impact the performance of the quality metrics. For these three models, FR metrics based either only on color features or both on geometry and color features usually achieve higher performance, with image-based metrics also having a very high correlation. For StMichael, correlation coefficients are somehow lower, although this same set of metrics continues to correlate well with human perception. However, for the sparser point clouds Boxer and House_without_roof, the performance of point cloud metrics drops, with image-based metrics achieving the highest correlation coefficients. This effect may be due to the way that different codecs handle the sparsity, since G-PCC always reduces the number of points at the same time that V-PCC and JPEG Pleno can dramatically increase their density as discussed in [1]. Since the point size is adapted in the rendering framework to produce watertight surfaces, the impact of these different effects on subjective perception is probably easier to capture directly in the projections than in the point domain. It is also observed that generally, geometry-only metrics perform worse than other FR metrics. While this effect is expected due to their lack of awareness about color distortion, this distinction is clearer in the evaluation of cross-codec generalization rather than cross-content generalization.
@@ V Conclusions
In this paper, a comprehensive set of objective quality metrics is assessed to obtain better insights on their correlation with subjective scores obtained in a point cloud quality experiment including compression artifacts generated by G-PCC, V-PCC, and JPEG Pleno. The evaluation is first conducted for the entire dataset where image-based metrics FSIM and VMAF were able to achieve the highest performance, with point cloud quality metrics PCQM, MS-PointSSIM, and GraphSIM coming next. The metrics were also separately assessed for cross-content and cross-codec generalization, showing that higher correlation coefficients can usually be achieved across different types of artifacts for the same point cloud rather than the other way around. A drop in performance is observed when only sparser point clouds are considered, suggesting that current point cloud metrics are better suited to estimate visual quality for solid models. Future work may consider the insights provided in this study to develop more robust quality metrics, especially given the importance of quality predictors for the development and improvement of point cloud compression solutions.

### v2

@@ IV Results and discussion
Although the performance indexes computed across the entire dataset provide a good indication of both cross-content and cross-codec generalization, there are scenarios where requirements may be lower. For instance, when a metric is used to estimate the quality of point clouds compressed with a known codec, only cross-content generalization could be desired, which can be better evaluated by the six right-most columns of Table I. It is observed that correlation values with G-PCC are usually higher than for the other codecs, with many metrics achieving PLCC higher than 0.9 and significantly better performance than for the entire dataset. A possible explanation is the simplicity of octree-based geometric artifacts, as stronger compression leads to fewer points in the decoded point cloud and sparser distribution. GraphSIM achieves the best performance in this case, followed by MS-GraphSIM and image-based VMAF.
However, several metrics were not able to maintain high performance for V-PCC, where sharp drops in performance are observed for both point-to-distribution variants, Histogram Y and YUV, MS-GraphSIM, GraphSIM, and PointSSIM. For the last two metrics, it can be observed in Figure 1 that stimuli generated from Boxer and House_without_roof with low objective scores received a high MOS. As it can be seen in Figure 2, V-PCC is able to keep relatively good quality even at the lowest rate, resulting in a high MOS value. However, the zoom on the hand of the model shows that the codec increases the point density, which cannot be perceived by subjects in the renderings displayed in the experiment but can affect the metrics score. Predictors computing features over neighborhoods defined with k-Nearest Neighbors such as GraphSIM and PointSSIM can be particularly affected since the neighborhoods will span over regions with smaller sizes for the denser model, potentially changing the distribution of the attributes. Image-based FSIM and PCQM achieve the highest performance, followed by MS-PointSSIM in terms of PLCC and MM-PCQA according to the SROCC. For JPEG Pleno, the maximum correlation values are in between G-PCC and V-PCC, with GraphSIM and LogP2D-JGY achieving the highest correlation. The performance of many metrics is more similar for JPEG Pleno and G-PCC than for V-PCC. For instance, Y PSNR and YUV PSNR are among the best-performing color-based metrics for V-PCC while the opposite is true for the other codecs. Likewise, several metrics that fail for V-PCC with PLCC under 0.4 have much better behavior for JPEG Pleno and G-PCC.
(a) Reference
(b) V-PCC R1
Refer to caption Refer to caption
MOS=4.75 PMOS: PCQM=4.31 GraphSIM=3.55 PointSSIM=2.99
Figure 2: Original Boxer and compressed with V-PCC at P1 and R1 displayed side-by-side, with a zoom on the hand region on the right. MOS value, as well as PMOS as given by Equation 1 for different metrics are also reported.
@@
The recent rise in research of point cloud compression techniques providing high subjective quality highlighted the need for metrics with high cross-codec generalization ability since the performance of new codecs is evaluated by comparison of its rate-distortion curve to anchor coding methods separately for different point clouds. Table II illustrates the PLCC and SROCC computed separately for each point cloud, which are indicators of cross-codec generalization. Most objective metrics achieve higher correlation when compared to previous evaluations and the highest coefficients are achieved for Bouquet, Soldier, and Thaidancer, which are solid point clouds, i.e. with very high point density. The higher voxelization precision used for Thaidancer of 12 bits does not seem to negatively impact the performance of the quality metrics compared to the other two point clouds where 10 bits were used. For these three models, FR metrics based either only on color features or both on geometry and color features usually achieve higher performance, with image-based metrics also having a very high correlation. For StMichael, correlation coefficients are somehow lower, although this same set of metrics continues to correlate well with human perception.
However, for the sparser point clouds Boxer and House_without_roof, the performance of point cloud metrics drops, with image-based metrics achieving the highest correlation coefficients. This effect is likely due to the way that different codecs, especially V-PCC, handle the sparsity of point clouds, as previously discussed and illustrated in Figure 2. Since the point size is adapted in the rendering framework to produce watertight surfaces, the impact of these different effects on subjective perception is probably easier to capture directly in the projections than in the point domain. It is also observed that generally, geometry-only metrics perform worse than other FR metrics. While this effect is expected due to their lack of awareness about color distortion, the performance penalty for even the best geometry-based predictor is observed to be higher for cross-codec generalization than for cross-content generalization.
@@ V Conclusions
In this paper, a comprehensive set of objective quality metrics is assessed to obtain better insights into their correlation with subjective scores obtained in a point cloud quality experiment including compression artifacts generated by JPEG Pleno, G-PCC, and V-PCC. The evaluation is first conducted for the entire dataset where image-based metrics FSIM and VMAF were able to achieve the highest performance, with point cloud quality metrics PCQM, MS-PointSSIM, and GraphSIM coming next. The metrics were also separately assessed for cross-content and cross-codec generalization, showing that higher correlation coefficients can usually be achieved across different types of artifacts than across different point clouds. A drop in performance is observed when only sparser point clouds are considered, suggesting that current point cloud metrics are better suited to estimate visual quality for solid models. Future work may consider the insights provided in this study to develop more robust quality metrics, especially given the importance of quality predictors for the development and improvement of point cloud compression solutions.